### v1

# Uncertainty-Aware Adaptation of Large Language Models for Protein-Protein Interaction Analysis

Abstract
Identification of protein-protein interactions (PPIs) helps derive cellular mechanistic understanding, particularly in the context of complex conditions such as neurodegenerative disorders, metabolic syndromes, and cancer. Large Language Models (LLMs) have demonstrated remarkable potential in predicting protein structures and interactions via automated mining of vast biomedical literature; yet their inherent uncertainty remains a key challenge for deriving reproducible findings, critical for biomedical applications. In this study, we present an uncertainty-aware adaptation of LLMs for PPI analysis, leveraging fine-tuned LLaMA-3 and BioMedGPT models. To enhance prediction reliability, we integrate LoRA ensembles and Bayesian LoRA models for uncertainty quantification (UQ), ensuring confidence-calibrated insights into protein behavior. Our approach achieves competitive performance in PPI identification across diverse disease contexts while addressing model uncertainty, thereby enhancing trustworthiness and reproducibility in computational biology. These findings underscore the potential of uncertainty-aware LLM adaptation for advancing precision medicine and biomedical research.
Index Terms:
Large Language Model (LLM), Low Rank Adaptation (LoRA), Uncertainty Quantification (UQ), Bayesian Inference, Deep Ensemble, Protein-Protein Interaction (PPI).
I Introduction
Protein–protein interactions (PPIs) form the molecular foundation of cellular function, orchestrating everything from gene regulation and signal transduction to metabolic processes and immune response. The intricate network of these interactions, often termed the interactome, represents one of the most complex and dynamic systems in biology [1]. Understanding this complex PPI network is particularly crucial in disease contexts, where aberrant protein interactions can lead to pathological states. Alterations in PPI networks have been implicated in numerous diseases, affecting fundamental cellular processes such as protein homeostasis, cell cycle regulation, and metabolic control. These disease-associated changes in the interactome can manifest through various mechanisms, from disrupted protein complex formation to altered signaling cascades, ultimately contributing to disease progression and severity. Elucidating these interaction networks is therefore essential for understanding disease mechanisms and developing therapeutic strategies [2].
Traditional experimental methods for identifying PPIs, such as yeast two-hybrid screening and co-immunoprecipitation, have generated vast amounts of validated interaction data. These experimentally determined interactions have been systematically collected in comprehensive databases such as STRING [3], BioGRID [4], and IntAct [5], creating valuable resources for the research community. While experimental methods remain the gold standard for PPI validation, their labor-intensive and time-consuming nature has motivated the development of computational approaches to predict novel protein interactions.
Early computational methods for PPI prediction primarily relied on sequence-based evolutionary patterns across species [6, 7, 8]. As the field advanced, machine learning (ML) approaches emerged, offering new ways to integrate multiple types of biological data [9]. These included convolutional neural networks (CNNs) for analyzing protein sequence patterns, recurrent neural networks (RNNs) for capturing sequential dependencies, and graph neural networks (GNNs) for modeling the topology of protein interaction networks [10].
The advancements of large language models (LLMs) [11, 12, 13, 14, 15, 16, 17] have transformed multiple scientific domains through their unprecedented capabilities in understanding complex patterns and relationships in text data [18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. Building on their powerful language-processing capabilities, biology-specific models have been developed to tackle diverse tasks: ProtBERT [28] and ESM3 [29] focus on protein sequence analysis, while BioGPT [19] and BioMedGPT [30] have shown promise in extracting biological knowledge from scientific literature. Recently, [31] has investigated PPI prediction using LLM-based approaches, demonstrating the potential of these models in specialized biomedical tasks. However, a critical challenge remains: such models often produce overly confident predictions, especially when trained on limited data, posing serious risks in high-stakes biomedical applications [32, 33, 34, 35]. Uncertainty-aware adaptation of LLMs is particularly critical in biomedical applications where miscalibrated confidence levels could lead to erroneous conclusions about disease mechanisms or therapeutic targets. By incorporating principled uncertainty estimation and confidence calibration techniques, we can ensure that LLM-driven predictions are not only powerful but also reliable for guiding biomedical discoveries.
Figure 1: Illustration of our uncertainty-aware low-rank adaptation approach for pre-trained LLMs in protein-protein interaction prediction.
Within the broader machine learning community, uncertainty quantification has long been a challenge due to the high-dimensional parameter spaces of deep learning models. Traditional Bayesian methods relying on Markov chain Monte Carlo sampling often become intractable at scale [36, 37]. Consequently, approximate Bayesian methods such as variational inference [38], sparse learning [39, 40], and dimension reduction [41] have been explored. Alternately, ensemble-based methods like deep ensembles [42], SWAG [43], and SeBayS [44] also provide principled strategies for capturing predictive variability. On the other hand, parameter-efficient fine-tuning techniques—particularly Low-Rank Adaptation (LoRA) [45]—makes it more tractable to adapt large models for specific downstream tasks. To this end, novel UQ-aware fine-tuning approaches have emerged that combine LoRA with Bayesian or ensemble-based ideas, such as Bayesian LoRA [46, 47, 48, 49] and LoRA ensembles [50, 51].
In this study, we specifically focus on LLaMA-3 [17] and BioMedGPT [30] as our primary LLM frameworks. We integrate LoRA-based fine-tuning with uncertainty-aware techniques to improve disease-specific PPI prediction. More concretely, we adopt Bayesian LoRA and LoRA ensemble methods to mitigate overconfidence and capture richer predictive variability. By leveraging the language-like structure of protein sequences, our approach naturally models intricate dependencies while generating well-calibrated estimates. To this end, our contributions include:
  1. 1.
    Low rank adaptation-based fine-tuning of LLaMA-3-8B and BioMedGPT-LM-7B, comparing standard LoRA fine-tuning with Bayesian LoRA and LoRA ensembles to tackle disease-focused PPI prediction.
  2. 2.
    Uncertainty quantification (UQ) integration to assess the reliability and robustness of PPI predictions.
  3. 3.
    Comprehensive uncertainty-aware evaluation across disease-specific protein interaction networks, specifically those relevant to neurodegenerative disorders, metabolic diseases, and cancer.
Overall, our results confirm that incorporating UQ strategies not only enhances PPI prediction accuracy but also yields better-calibrated confidence measures—critical for drawing robust conclusions in biomedical research. By advancing uncertainty-aware methods in LLM-based modeling, we lay the groundwork for safer, more reliable, and more informative computational tools in precision medicine.
II Preliminaries
Throughout the paper, all vectors and matrices are denoted by bold lowercase $(\bm{l})$ and uppercase letters $(\bm{L})$ respectively.
II-A Low-Rank Adaptation
To adapt a pre-trained language model to downstream tasks, The authors of [45] introduced LoRA, a parameter-efficient fine-tuning approach. Assuming that weight changes exhibit a low intrinsic rank, LoRA optimizes rank decomposition matrices while keeping the pre-trained weights frozen.
Specifically, given that the weight update has a low-rank structure, the adapted forward pass is expressed as:
$\bm{h}=(\bm{W}_{0}+\Delta\bm{W})\bm{a}=(\bm{W}_{0}+\bm{B}\bm{A})\bm{a}.$
Here $\bm{a}$ and $\bm{h}$ represent the input and output vectors, respectively, of a large frozen pre-trained weight matrix $\bm{W}_{0}\in\mathbb{R}^{d_{1}\times d_{2}}$. The matrices $\bm{B}\in\mathbb{R}^{d_{1}\times r}$ and $\bm{A}\in\mathbb{R}^{r\times d_{2}}$ contain trainable parameters, with $r\ll\min(d_{1},d_{2})$. This reduction in the number of parameters, allows LoRA to provide efficient fine-tuning with a decrease in storage requirements. We adopt this model as a baseline in our experiments.
II-B Bayesian model formulation
Let $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1,\cdots,N}$ be a training dataset with $N$ i.i.d. samples, where $\bm{x}$ represents the input samples and $y$ represents the output samples. Bayesian inference captures model uncertainty by inferring a probability distribution over model parameters, $\bm{\theta}=(\theta_{1},\cdots,\theta_{T})\in\mathbb{R}^{T}$, instead of learning a single deterministic model–$p(y|\bm{x},\bm{\theta})$. The posterior distribution follows Bayes’ rule: $p(\bm{\theta}|\mathcal{D})\propto p(\mathcal{D}|\bm{\theta})p(\bm{\theta})$, where $p(\mathcal{D}|\bm{\theta})$ is the model likelihood and $p(\bm{\theta})$ is the prior distribution. To make predictions for a new input $\bm{x}_{\rm new}$, Bayesian model averaging (BMA) is applied using the posterior distribution $p(\bm{\theta}|\mathcal{D})$ as follows:
[…]
$\approx\frac{1}{B}\sum_{b=1}^{B}p(y_{\rm new}|\bm{x}_{\rm new},\bm{\theta}_{b}),\enskip\bm{\theta}_{b}\sim p(\bm{\theta}|\mathcal{D}).$
This approach improves generalizability and model calibration by incorporating parameter uncertainty into predictions.
III Methodology
III-A LoRA Ensemble
We employ an ensemble of LoRA models – LoRA Ensemble [50, 51] as an efficient strategy for uncertainty quantification in LLMs. Traditional deep ensembles yield better predictive performance and uncertainty estimation by training multiple models independently, but applying this directly to LLMs is often infeasible due to high memory and computational costs.
To circumvent these issues, each LoRA Ensemble member fine-tunes the same pre-trained backbone $\bm{W}_{0}$ with a low-rank trainable modification $\Delta\bm{W}_{m}=\bm{B}_{m}\bm{A}_{m}$, where $\bm{B}_{m}\in\mathbb{R}^{d_{1}\times r}$ and $\bm{A}_{m}\in\mathbb{R}^{r\times d_{2}}$ have significantly fewer parameters than the full model, $r_{m}\ll\min(d_{1},d_{2})$. These adapters are trained independently and in parallel, ensuring diverse solutions–$\{\bm{W}_{1},\bm{W}_{2},\dots,\bm{W}_{M}\}$. The ensemble prediction is computed by averaging outputs across $M$ ensemble members. For a given input $\bm{x}_{\rm new}$, if $y^{m}_{\rm new}$ represents the prediction from the $m$-th ensemble member, the final ensemble output (for continuous outcomes) is given by:
$p_{\rm ens}(y_{\rm new}|\bm{x}_{\rm new})=\frac{1}{M}\sum_{m=1}^{M}p(y^{m}_{\rm new}|\bm{x}_{\rm new},\bm{W}_{m}).$
This approach retains the benefits of ensembling–improved accuracy, calibration, and robustness–while preserving efficiency by reusing the frozen backbone and only training lightweight LoRA adapters.
III-B Bayesian Low-Rank Adaptation
Despite the availability of scalable posterior inference methods like variational inference [38], a fully Bayesian treatment of LLMs remains computationally prohibitive. Instead, limiting Bayesian inference to LoRA parameters offers a more tractable means of capturing uncertainty in model predictions. However, even Markov chain Monte Carlo approaches can become excessively costly for inferring posteriors over the millions of LoRA parameters involved in large-scale models. As a practical compromise, Bayesian LoRA [46] employs the Laplace approximation to estimate the posterior over these low-rank parameters, centered around their maximum a posteriori (MAP) estimate together with covariance equaling the Fisher information matrix [52].
To this end, let $\bm{\theta}$ denote the trainable LoRA parameters with a prior distribution of $\mathcal{N}(\bm{0},\lambda^{-1}\bm{I})$. The Laplace approximation first calculates MAP estimate which is equivalent to maximizing the log-joint, $\log{\rm P}(\bm{y},\bm{X},\bm{\theta})$
$\bm{\theta}_{\rm MAP}$ $=\underset{\bm{\theta}}{\arg\!\max}\log{\rm P}(\bm{y},\bm{X},\bm{\theta})$
$=\underset{\bm{\theta}}{\arg\!\max}\log{\rm P}(\bm{y}|\bm{X},\bm{\theta})+\log{\rm P}(\bm{\theta})$
$=\underset{\bm{\theta}}{\arg\!\max}\log{\rm P}(\bm{y}|\bm{X},\bm{\theta})+\frac{\lambda}{2}||\bm{\theta}||^{2}_{2}+{\rm const}$
where $\bm{X}$ represents the model inputs. The term associated with log of the prior distribution provides us $L_{2}$-regularization on the trainable parameters. We can incorporate this in frequentist model training via weight decay term with $\lambda/2$ strength. As a result, parameters from any previously trained model that used a reasonable weight decay setting (for example, via AdamW with its weight decay) can be directly reused.
Next, to obtain an approximate posterior around $\bm{\theta}_{\rm MAP}$, Laplace method proceeds with a second-order Taylor expansion of the log-joint $\mathcal{L}(\mathcal{D},\bm{\theta})=\log p(\bm{y},\bm{X},\bm{\theta})$ around $\bm{\theta}_{\rm MAP}$. Hence, by ignoring the higher-order terms, this yields
$\mathcal{L}(\mathcal{D},\bm{\theta})\approx\mathcal{L}(\mathcal{D},\bm{\theta}_{\rm MAP})+\frac{1}{2}(\bm{\theta}-\bm{\theta}_{\rm MAP})^{\top}\bm{H}\,(\bm{\theta}-\bm{\theta}_{\rm MAP}),$
where the first-order term zeros out due to the zero gradient at $\bm{\theta}_{\rm MAP}$ and $\bm{H}$ is the Hessian of the log-joint at $\bm{\theta}_{\rm MAP}$, $\nabla^{2}_{\bm{\theta}}\mathcal{L}(\mathcal{D},\bm{\theta})|_{\bm{\theta}_{\rm MAP}}$. Under this quadratic approximation,
$p(\bm{\theta}\mid\mathcal{D})\approx\mathcal{N}\bigl{(}\bm{\theta}|\bm{\theta}_{\rm MAP},\,\bm{H}^{-1}\bigr{)}.$ (1)
Hence, Laplace approximation turns out to be post-hoc Bayesian inference method which requires the additional step of computing the $H^{-1}$ matrix at $\bm{\theta}_{\rm MAP}$. In practice, computing the full Hessian $\bm{H}$ can be expensive, especially for large models due to quadratic complexity with respect to the number of model parameters. We use the positive semi-definite Fisher information matrix to circumvent the issue of the potentially indefinite Hessian, which arises when local convexity conditions fail to hold in large machine learning models. Accordingly, the Fisher information is defined by
$\bm{F}(\bm{\theta})=\sum_{n}^{N}\mathbb{E}_{\hat{y}\sim{\rm P}(y|f_{\bm{\theta}}(\bm{x}_{n}))}\bigl{[}\bm{G}\bm{G}^{\top}\bigr{]}$
where $G=\nabla_{\bm{\theta}}{\rm P}(\hat{y}|f_{\bm{\theta}}(\bm{x}_{n}))$ represents the gradient and the expectation above is over the model’s output distribution. Next, in order to estimate the Fisher information in a manner that is both tractable and memory-efficient, we employ a Kronecker-Factored Approximate Curvature (K-FAC) approach similar to [46]. In K-FAC, we treat Fisher as a block-diagonal matrix for each linear layer and factorize each block into two smaller matrices. For the $l$-th linear layer, we compute Fisher block $F_{l}$ using that layer’s input activations $\bm{a}_{l-1}$ and log-likelihood gradients with respect to layer’s pre-activation output $s_{l}$ denoted by $\bm{G}_{\bm{s}_{l}}=\nabla_{\bm{s}_{l}}\log{\rm P}(\bm{y}|\bm{X},\bm{\theta})$. Hence the expression is
$\bm{F}_{l}=\sum_{n=1}^{N}\mathbb{E}_{{\rm P}(y|f_{\bm{\theta}}(\bm{x}_{n}))}\bigl{[}\bm{a}_{l-1}\bm{a}_{l-1}^{\top}\bigr{]}\otimes\mathbb{E}_{{\rm P}(y|f_{\bm{\theta}}(\bm{x}_{n}))}\bigl{[}\bm{G}_{\bm{s}_{l}}\bm{G}_{\bm{s}_{l}}^{\top}\bigr{]}$ (2)
This approach avoids storing the full, dense Hessian, thereby reducing computational overhead. By applying K-FAC to the LoRA parameters, we maintain a compact representation of uncertainty while keeping the overhead similar to standard training. However, in Equation (2), the first expectation grows with the square of the layer’s input width, while the second grows with the square of the output width. Because LoRA adapters alternate between wide-input-narrow-output configuration and vice versa, one of these expectations can become especially large. To address this, we use an incremental SVD to factorize the large matrix into two new low-rank factors thereby saving memory. Further mathematical details are provided in Appendix E of [46].
Once we infer the approximate posterior which is Gaussian as per Equation 1, we can linearize the model predictions around the MAP estimate $\bm{\theta}_{\mathrm{MAP}}$ [53]. For a test input $\mathbf{x}_{\rm new}$,
$f_{\bm{\theta}}(\bm{x}_{\rm new})\approx f_{\bm{\theta}_{\rm MAP}}(\bm{x}_{\rm new})+\nabla_{\bm{\theta}}f_{\bm{\theta}}(\bm{x}_{\rm new})\bigl{|}_{\bm{\theta}_{\rm MAP}}^{\top}\bigl{(}\bm{\theta}-\bm{\theta}_{\rm MAP}\bigr{)}.$
Because this expression is linear in $\bm{\theta}$, integrating out the Gaussian posterior over $\bm{\theta}$ yields a Gaussian predictive distribution for the logits:
$f_{\bm{\theta}}(\bm{x}_{\rm new})$ $\sim\mathcal{N}\bigl{(}\bm{y}|f_{\bm{\theta}_{\rm MAP}}(\bm{x}_{\rm new}),\Lambda\bigr{)},$
wh ere $\text{ere }\Lambda=\nabla_{\bm{\theta}}f_{\bm{\theta}_{\rm MAP}}(\bm{x}_{\rm new})^{\top}H^{-1}\nabla_{\bm{\theta}}f_{\bm{\theta}_{\rm MAP}}(\bm{x}_{\rm new}).$
Finally to efficiently sample from this predictive posterior, we use the Cholesky decomposition of $\Lambda=\mathbf{L}\mathbf{L}^{\top}$. Then,
$\hat{\bm{y}}=f_{\bm{\theta}}(\bm{x}_{\rm new})=f_{\bm{\theta}_{\rm MAP}}(\bm{x}_{\rm new})+\bm{L}\bm{z},\quad\bm{z}\sim\mathcal{N}\bigl{(}\bm{0},\bm{I}\bigr{)}.$
This linearized predictive step, combined with a Gaussian approximate posterior, yields efficient uncertainty estimates in Bayesian LoRA approach for downstream tasks.
TABLE I: ND-PPI Prediction: The best results among all compared methods for a given LLM pre-trained model are highlighted in bold. All metrics are reported as means with standard deviations in subscript, based on three independent runs.
LLM Model Methods Acc ($\uparrow$) NLL ($\downarrow$) ECE ($\downarrow$) Spec. ($\uparrow$) Prec. ($\uparrow$) F1 ($\uparrow$) MCC ($\uparrow$) AUROC ($\uparrow$)
Llama-3 Single LoRA 86.510.54 0.3620.036 0.0950.011 0.9660.005 0.8810.005 0.8630.006 0.7450.011 0.9530.004
LoRA Ensemble 88.700.62 0.3020.002 0.0880.005 0.9730.001 0.8990.004 0.8860.006 0.7850.011 0.9640.001
Bayesian LoRA 86.510.20 0.3170.003 0.0520.021 0.8480.033 0.8660.002 0.8650.002 0.7320.003 0.9440.004
BioMedGPT Single LoRA 85.442.16 0.5390.053 0.1190.025 0.9630.016 0.8740.011 0.8520.023 0.7260.034 0.9440.005
LoRA Ensemble 88.001.19 0.3630.049 0.0870.022 0.9650.008 0.8920.007 0.8790.013 0.7710.019 0.9560.001
Bayesian LoRA 86.820.60 0.3200.012 0.0330.007 0.8690.015 0.8680.006 0.8680.006 0.7370.012 0.9370.003
TABLE II: M-PPI Prediction: The best results among all compared methods for a given LLM pre-trained model are highlighted in bold. All metrics are reported as means with standard deviations in subscript, based on three independent runs.
LLM Model Methods Acc ($\uparrow$) NLL ($\downarrow$) ECE ($\downarrow$) Spec. ($\uparrow$) Prec. ($\uparrow$) F1 ($\uparrow$) MCC ($\uparrow$) AUROC ($\uparrow$)
Llama-3 Single LoRA 85.820.26 0.3980.016 0.0840.006 0.9080.036 0.8630.007 0.8580.002 0.7210.009 0.9370.002
LoRA Ensemble 87.450.16 0.3080.013 0.0510.010 0.9220.016 0.8780.003 0.8740.002 0.7520.004 0.9500.002
Bayesian LoRA 83.411.17 0.3740.005 0.0710.018 0.9320.038 0.8500.003 0.8320.013 0.6830.013 0.9250.004
BioMedGPT Single LoRA 83.680.54 0.5420.026 0.1130.009 0.7990.047 0.8400.002 0.8370.006 0.6780.007 0.9250.005
LoRA Ensemble 87.141.39 0.3540.028 0.0620.010 0.8640.036 0.8720.013 0.8710.014 0.7440.027 0.9410.005
Bayesian LoRA 83.290.57 0.3850.015 0.0370.018 0.8880.033 0.8380.003 0.8320.007 0.6710.007 0.9050.004
TABLE III: C-PPI Prediction: The best results among all compared methods for a given LLM pre-trained model are highlighted in bold. All metrics are reported as means with standard deviations in subscript, based on three independent runs.
LLM Model Methods Acc ($\uparrow$) NLL ($\downarrow$) ECE ($\downarrow$) Spec. ($\uparrow$) Prec. ($\uparrow$) F1 ($\uparrow$) MCC ($\uparrow$) AUROC ($\uparrow$)
Llama-3 Single LoRA 96.620.62 0.0940.011 0.0330.002 0.9730.016 0.9670.007 0.9660.006 0.9320.013 0.9960.002
LoRA Ensemble 97.860.00 0.0660.005 0.0290.005 0.9800.000 0.9790.000 0.9790.000 0.9570.000 0.9970.000
Bayesian LoRA 96.971.24 0.0850.020 0.0270.002 0.9630.012 0.9700.012 0.9700.012 0.9400.025 0.9960.002
BioMedGPT Single LoRA 97.680.82 0.0590.011 0.0250.011 0.9760.006 0.9770.008 0.9770.008 0.9540.016 0.9980.001
LoRA Ensemble 98.400.54 0.0520.000 0.0210.002 0.9800.000 0.9840.005 0.9840.005 0.9680.011 0.9980.000
Bayesian LoRA 98.400.54 0.0640.005 0.0310.001 0.9760.006 0.9840.005 0.9840.005 0.9680.011 0.9980.001
IV Experimental Results
In this section, we assess the performance of two uncertainty-aware LoRA adaptations—LoRA Ensemble and Bayesian LoRA—applied to LLaMA-3-8B and BioMedGPT-LM-7B models on publicly available protein-protein interaction datasets. As a baseline, we include a single LoRA model trained in a deterministic manner. All LoRA-based approaches were implemented using the PEFT library [54], with each configuration run three times using different random seeds. We evaluate model performance and robustness by accuracy (Acc), negative log-likelihood (NLL), and expected calibration error (ECE) on the test sets. Additional details on the NLL and ECE metrics can be found in Appendix V-A. Furthermore, we report Matthews Correlation Coefficient (MCC), specificity (Spec.), precision (Prec.), F1-score, and Area under Receiver Operating Characteristic curve (AUROC) over test sets for a comprehensive view of predictive capabilities. Final metrics are summarized by the mean and standard deviation across three independent runs.
PPI Datasets. The datasets analyzed here explore PPIs related to various diseases, providing valuable insights into their underlying mechanisms. The Neurodegenerative diseases PPI (ND-PPI) dataset, sourced from the study [55], focuses on neurodegenerative diseases and examines a network of 820 proteins forming 11,762 interactions, evenly split between positive and negative pairs. The dataset is structured to assess whether specific protein pairs interact in the presence of neurodegenerative conditions. Similarly, the metabolic disorders PPI (M-PPI) dataset, also from [55], investigates metabolic disorders, encompassing 1,063 proteins and a total of 10,262 interactions. The cancer PPI (C-PPI) dataset, derived from the study [56], consisted of 933 positive and 1,308 negative interactions. To ensure balanced representation, this dataset was curated to create an equal-sized collection of 1,866 total interactions. These datasets have been evaluated by prompting models to assess whether proteins interact in the corresponding conditions and collectively contribute to advancing computational models for predicting PPIs across different disease contexts, enhancing our understanding of disease-specific interaction networks. These tasks are formalized into binary (True/False) classification problems as illustrated in Fig. 1. Furthermore, each dataset is divided into $80\%$ for training and $20\%$ for testing, with all models evaluated on the fixed test set in each PPI prediction task. We refer readers to [31], for additional details and exploratory analyses of these datasets.
Implementation Details. In all experiments, we construct a LoRA ensemble using three individually fine-tuned LoRA learners. The LoRA matrices $\bm{B}$ are initialized to zero, while the entries of $\bm{A}$ follow a Kaiming Uniform initialization [57]. Optimization is performed using the AdamW optimizer with a learning rate of $1\times 10^{-4}$, default hyperparameters, and a total of four training epochs. The batch size is set to $4$ for the ND-PPI and M-PPI cases and $16$ for the C-PPI case, following [31]. For Bayesian LoRA, the prior precision $\lambda$ is fixed at $0.1$. Lastly, LoRA is applied to the queries, values, and output layer across all methods, with specific hyperparameters set to $r=16$, $\alpha=32$, a dropout rate of $0.05$, and a maximum sequence length of $50$.
Results. The results for ND-PPI, M-PPI, and C-PPI tasks are summarized in Tables I, II, and III, respectively.
In the ND-PPI prediction task (Table I), we demonstrate that the LoRA ensemble achieves the highest predictive accuracy among all models in both LLM settings and has the lowest NLL in the LLaMA-3 fine-tuning case. Conversely, Bayesian LoRA demonstrates the best calibration in both scenarios, exhibiting the lowest ECE and achieving the lowest NLL in the BioMedGPT fine-tuning case. Lastly, the LoRA ensemble reports the highest values for specificity, precision, F1-score, MCC, and AUROC among all models. In the M-PPI prediction task (Table II), we show that the LoRA ensemble achieves the highest predictive accuracy and lowest NLL in both LLM scenarios, while also attaining the lowest ECE in the LLaMA-3 case. Conversely, Bayesian LoRA achieves the best calibration in the BioMedGPT case and the highest specificity in both scenarios. Finally, the LoRA ensemble outperforms all the models by achieving best precision, F1-score, MCC, and AUROC values.
In the C-PPI prediction task (Table III), we demonstrate that the LoRA ensemble once again achieves the highest predictive accuracy and lowest NLL in both settings, while also attaining the lowest ECE in the BioMedGPT scenario. Bayesian LoRA matches the best predictive accuracy in the BioMedGPT case and achieves the lowest ECE in the LLaMA-3 case. In the LLaMA-3 setting, the LoRA ensemble reports the highest values for specificity, precision, F1-score, MCC, and AUROC among all models. Additionally, it achieves the best specificity in the BioMedGPT case. Notably, both Bayesian LoRA and the LoRA ensemble attain the best precision, F1-score, and MCC values in the BioMedGPT case. Lastly, all three models yield identical AUROC values in the BioMedGPT case.
V Conclusion and Discussion
In this study, we presented a novel uncertainty-aware adaptation of LLMs approach for predicting protein-protein interactions across multiple disease contexts. Leveraging fine-tuned LLaMA-3 and BioMedGPT models with LoRA ensemble and Bayesian LoRA, our approach consistently improved prediction accuracy, reliability, and robustness, as confirmed by comprehensive metrics such as negative log-likelihood and calibration error. LoRA ensembles excelled at achieving higher accuracy and reliable uncertainty estimates, while Bayesian LoRA provided well-calibrated predictions. Together, they demonstrated robustness in neurodegenerative, metabolic, and cancer-related PPI tasks. These findings underscore the benefits of incorporating principled uncertainty quantification into parameter-efficient fine-tuning for LLMs. Future work will explore more advanced LLM uncertainty quantification methods and apply this methodology to broader biomedical applications. Potential directions include elucidating disease mechanisms by predicting disrupted protein interactions and, consequently, advance computational protein target discovery for therapeutic design.
APPENDIX
V-A Robustness & Predictive Uncertainty Evaluation Metrics
To assess model robustness and predictive uncertainty, we use Negative Log-Likelihood (NLL) and Expected Calibration Error (ECE). NLL evaluates how confidently a model predicts the correct labels. Given a test dataset $\{\bm{x}_{i},y_{i}\}_{i=1}^{N}$, NLL is computed as:
$\text{NLL}=\frac{1}{N}\sum_{i=1}^{N}-\log{\rm P}_{\bm{\theta}}(y_{n}).$
A lower NLL indicates better confidence calibration, as overconfident incorrect predictions increase this value. On the other hand, ECE measures how well predicted confidence aligns with actual accuracy. Predictions are grouped into bins based on confidence, and ECE is calculated as:
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|.$
Here, $\text{acc}(B_{m})$ and $\text{conf}(B_{m})$ represent the average accuracy and confidence within bin $B_{m}$, respectively:
$\text{acc}(B_{m})$ $=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\mathbf{1}(\hat{y}_{i}=y_{i}),$
$\quad\text{conf}(B_{m})$ $=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}P(\hat{y}_{i}),$
where $|B_{m}|$ is the number of samples in bin $m$. Across all experiments, we set $|B_{m}|=15$.